OpenStudy (desmarie):

Find the area of the region between a regular hexagon with sides of 6" and its inscribed circle.

OpenStudy (desmarie):

@imqwerty

imqwerty (imqwerty):

area of regular hexagon is given by this formula-\(\large \frac{3 \sqrt{3}}{2}a^2\) where \(a\) is the side length of the hexagon finding the radius of circle- here i have not drawn the circle tho we can use trigonometric ratios to find the radius hint: \(tan(60)= \large \frac{radius}{a/2}\) once you find this radius the area of circle will be this- \(\pi r^2\) and then all you need to do is subtract the area of circle from the area of hexagon to get your answer :)